Circular No. 6324

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 M. J. Stark, University of Maryland and the Laboratory for High

Energy Astrophysics, Goddard Space Flight Center (LHEA/GSFC); A. Baykal, National Research Council and LHEA/GSFC; T. Strohmayer, Universities Space Research Association and LHEA/GSFC; and J. H. Swank, LHEA/GSFC, communicate: “We have detected, with the Proportional Counter Array (PCA) onboard the X-ray Timing Explorer (XTE), glitches in the arrival times of x-ray pulses from GRO J1744–28 that occur during the x-ray/γ-ray bursts seen from this source. Timing of the pulses before and after the burst events shows a clear advance in the pulse arrival times of ∼ 25 ms (5 percent of the 0.467-s period). The advance occurs within a time interval no longer than 25 or 30 s, after which the advance decays. The functional form of this decay is not yet precisely known, but the bestfit of an exponential function to it has a time constant of ∼ 500 s. This behavior is exhibited during all bursts so far observed from this source by XTE, and these pulse timing considerations have been used to identify additional bursts that occurred when the PCA was not operating. The flux from GRO J1744–28 is depressed slightly following a burst but recovers to its pre-burst level within a few hundred seconds, while the phase shift recovers in ∼ 1500 s. The pulse shapes before and after the bursts are very similar. Interpreted in terms of a fractional change in the angular velocity, the short time during which the advance occurs indicates such a change of at least 10^{−3}, whereas the largest such change seen in radio-pulsar glitches is 10^{−6}. If 10^{20} g of material with the angular momentum that is characteristic of the co-rotation radius (10^{8} cm) are accreted, the fraction of the star’s moment of inertia participating in the glitch is ∼ 10^{−6}, equivalent to the moment of inertia of a thin shell containing 10^{−6} of the neutron-star mass. This is much less than the fraction of the moment of inertia thought to participate in radio-pulsar glitches.”

COMET 22P/KOPFF Total magnitude estimates: Feb. 18.19 UT, 13.0 (H. Mikuž, Ljubl-

jana, Slovenia, 0.20-m reflector + CCD + V filter); 19.51, 12.4 (A. Hale, Cloudcroft, NM, 0.41-m reflector, visual); 23.46, 12.2 (Hale).

1996 February 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green